A machine does 1920 joule of work in 240 second. what is the power of the machine ?

Grade:11

3 Answers

varsha
6 years ago
power =work done/time taken
          = energy/time taken 
         = 1920/240
         =8 J/s
therefore power of that machine is 8 J/s
somi teez
6 years ago
hi,
power of anything is defined as rate of change of work done (at what rate you are doing a work)
P=W/t
  =1920 J/ 240 sec
  =8 J/sec
 =8 watt
Hence,work done is 8 watt
 
Indestructible Avinash
6 years ago
We know that power is equal to work done per unit time given work (w)=1920 time t =240p=1920/240 =8Therefore power=8watt